will has $2.05 in dimes and quarters.He has 4 more quarters than dimes.Find the number he has of each kind of coin
Answer by checkley77(12569)

D=Q+4
.25Q+.10D=2.05
.25Q+.10(Q+4)=2.05
.25Q+.10Q+.4=2.05
.35Q=2.05-.4
.35Q=1.65
Q=1.65/.35
Q=4.714 NUMBER OF QUARTERS.
D=4.714+4
D=8.714 DIMES NUMBER OF.
PROOF:
.25*4.714+.10*8.714=2.05
1.1785+.8714=2.05
205~205
CAN'T HAVE PARTIAL COINS. THERE MUST BE A MISTAKE IN THE COIN COUNT OR THE TOTAL AMOUNT OF THE COINS.